OpenStudy (anonymous):

Ivan rented a truck for one day. There was a base fee of $14.99 , and there was an additional charge of 97 cents for each mile driven. Ivan had to pay 186. 68 when he returned the truck. For how many miles did he drive the truck?

OpenStudy (anonymous):

total money= base fee + no of miles *0.97 , therefore 186.68 =14.99 + no of miles*0.97 , no of miles*0.97=186.68 -14.99 , no of miles*0.97=171.69 , no of miles =171.69/0.97=177 miles